A 2 lb rock is dropped from a bridge to the water that is 50 ft below. What is its kinetic energy at the moment it hits the water? Submit your answer in US units to the nearest tenth of a unit.
Answer by Fombitz(32388)

KE=(1/2)mv^2
v^2=2gh
So then
KE=(1/2)m(2gh)=mgh
The maximum KE is equal to the max. potential energy.
Plug in the values.